radomes.openskiesproject.org
shylacoplen1 shylacoplen1
  • 18-07-2016
  • Mathematics
Answered

fred randomly draws two cards from a standard deck of 52 cards he dose not replace the first card what is the probability there both jacks?

Verified Answer :

susie098
susie098 susie098
  • 18-07-2016

4/52 * 3/51 = 12/2652 = .452%

Answer Link
toporc
toporc toporc
  • 18-07-2016

The probability that the first card drawn is a jack is 4/52. If a jack was drawn first, there would be 3 jacks remaining in a total of 51 cards. Therefore the probability of drawing two jacks is given by:
\(P(2\ jacks)=\frac{4}{52}\times\frac{3}{51}=\frac{1}{221}\)

Answer Link

Related Questions

7. A student would like to show the difference in position between terrestrial planets and gas planets. Which type of planets are closer to the sun? Terrestri
I need the coordinates of each point under the given transformation
CTHIG Hale VI LIG VICU Nation15According to this quotation from its charter what isthe mission of the United NationsPlace the world under a single governmentImp
what is the anwser to 5x38
Symbols placed above vowels and some consonants in order to aid pronunciation are called marks diacritical marks accent points phonemic indicators
What is the difference between immigration and settler colonialism
What is the exact value of the expression 72 8 128 Simplify if possible8212283123
Math help plsss on these questions
Identify the subject in the following sentence Bob put a fly on his line
Why would a totalitarian government be considered to be more efficient at making decisions than a democratic one